# Geometrical properties of strained and twisted moiré heterostructures

### v1

@@ IV.2 Strained moiré patterns
As discussed in previous sections, the strain is a powerful tuning knob for manipulating the size and symmetry of the moiré patterns, much beyond what is possible via only a twist angle. In this section, we provide experimentally examples of nonhexagonal moiré geometries arising due to combinations of twist and strain. Here, the strain could be intentionally generated through well-developed strain techniques (Section IV.1), unintentionally induced during the sample fabrication, or appear due to the lattice relaxation. We will first illustrate a continuous approach to fine tune the twist and strain [kapfer2023programming], and then show particular observations of rectangular patterns [ou2025continuous], quasi-one-dimensional patterns [alden2013strain, bai2020excitons, mendoza2021strain, shabani2021deep], and giant atomic swirl deformations [mesple2023giant].
@@ IV.2.1 Continuous tuning moiré patterns
In the moiré systems, the geometric and electronic properties are often highly dependent on the moiré periodicity and the interlayer atomic registry. For example, the superconductivity in TBG is observed around a twist angle 1.1, but has a tolerance of only 0.1 [cao2021nematicity]. Therefore, a further optimization of the moiré electronic properties requires a technique to precisely and continuously tune the moiré pattern with twist angle and strain [kapfer2023programming, pantaleon2024designing].
Kapfer et al. develop a continuous approach to fine tune the twist angle and strain in a graphene–hBN heterostructure, based on the in-plane bending of monolayer ribbons through an AFM tip [kapfer2023programming]. Figure 13 illustrates the continuous tuning approach: a monolayer graphene (MLG) ribbon is bent by a few-layer graphite (FLG) manipulator shaped into a gear-like geometry and an AFM tip. The strain induced by bending the ribbon varies both in the longitudinal (x) and transverse (y) directions. The strain map in the bent ribbon follows some general tendencies: (i)  the strain changes linearly in both directions and has the maximum value at the end of the ribbon; (ii)  in the transverse direction, the ribbon has compressive and tensile strain on the inside and outside radius, respectively, which are separated by an unstrained neutral axis; (iii)  the maximum strain in the transverse and longitudinal directions are linearly dependent on the width $W$ and length $L$ of the ribbon, respectively. Therefore, this continuous tuning approach allows one to manipulate independently the twist angle and strain in 2D materials. Importantly, the bending process is robust and reversible.
@@ IV.2.2 Rectangular moiré patterns
Figure 14 shows a strain-induced moiré pattern manipulation in twisted bilayer WSe2 [ou2025continuous]. In this setup, a continuous and homogeneous uniaxial heterostrain is generated by bending a flexible substrate with different radius $R$ (see Sec. IV.1.1). The deformed moiré patterns were visualized by a piezoresponse force microscopy (PFM). As discussed in the previous subsection, in the bending substrate technique the strain is only transferred from the substrate to the bottom layer WSe2, generating a heterostrain in the moiré pattern. In the experiment, the substrate was mounted on several molds with different curvatures, which controlled the strain magnitude from 0% (flat) to 0.5% (radius curvature of 12.5 mm). As seen in Figures 14(b)-(f), this strain variation results in a continuous symmetry modulation of the moiré patterns, from triangular ($\beta_{R}=120^{\circ}$) to rectangular ($\beta_{R}=90^{\circ}$).
Figure 15: Strain induced 1D pattern in TBG. Non-local nano-photocurrent map of a TBG sample with $\theta<0.1^{\circ}$. In the photocurrent map, the bright spots highlight the AA sites. The AB and BA regions are indicated by orange and cyan dots, respectively. In the strained region, double domain walls (DDWs) are formed, separating regions of identical stacking configurations (AB or BA stackings). The green network overlaid on the photocurrent data corresponds to the prediction with one tuning parameter $\bar{\beta}$, indicating the ratio of DDW and single domain wall (SDW) line energies. Adapted under the terms of the $C$ BY license from Ref. [halbertal2021moire]. Copyright (2021) The author(s).
@@ IV.2.3 Quasi-1D pattern
As discussed in Section III.4.1, in strained moiré materials there can be a critical strain configuration at which the moiré collapse into a quasi-1D pattern. The critical condition can be, in principle, engineered by extenally induced-strain (Section IV.1). But quasi-1D pattern can also emerge most naturally by unintensional strains, either due to lattice relaxation or sample fabrication. For instance, Figure 15 shows clear rich features of quasi-1D moiré strips of relaxed DWs, for example, the shapes of domains, single domain walls (SDWs), and formation of double domain walls (DDWs) [halbertal2021moire]. A dimensionless parameter as the ratio of DDW and SDW line energies was defined to adress a competition between SDWs and DDWs. For $\bar{\beta}<2$, two SDWs attracted each other favoring the emergence of DDW segments. The SDWs and DDWs are also observed in some other moiré materials [halbertal2021moire, alden2013strain, halbertal2022unconventional].
Similar strain-induced quasi-1D patterns have been also observed in, for example, TBG strained by a substrate [mendoza2021strain], large-Angle Twisted Bilayer WTe2 [yang2025intrinsic], and exfoliated highly oriented pyrolytic graphite [boi2025quasi].
Figure 16: Strain induced giant swirl in bilayer graphene. STM current image ($V_{b}=-300$ mV, $I_{t}=250$ pA) for an intercalated surface of bilayer graphene on SiC, with a deformed moiré showing an (anti-) clockwise swriling feature in (blue) orange. NI indicate the non-intercalated regions. The inset is the STM image of the white dotted square region measued at $V_{b}=900$ mV, $I_{t}=250$ pA. The scal bar is 100 nm. Adapted under the terms of the $C$ BY license from Ref. [mesple2023giant]. Copyright (2023) The author(s).
@@ IV.2.4 Giant atomic swirl
As noted in Section III.6, relaxation-induced giant atomic swirls can emerge in the domain wall networks of moiré materials [mesple2023giant, shi2025spontaneous]. Figure 16 shows a moiré pattern formed in bilayer graphene with only biaxial heterostrain, which was produced by hydrogen intercalation of graphene on SiC [mesple2023giant]. The strain was only imposed to the bottom graphene layer by the SiC reconstruction. Due to the presence of lattice reconstruction in the moiré pattern, domain walls connecting AA stacking formed a triangular network of AB/BA stackings. Notably, giant atomic swirls appeared spontaneously around the AA regions. The neighboring twirl chiralities tend to anti-align (blue and orange arrows). Similar flower-shaped domain walls were already reported in trilayer epitaxial graphene on silicon carbide [lalmi2014flower].
@@ VI Acknowledgements
F.E. acknowledges support funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101210351. F.G acknowledges the support from the Department of Education of the Basque Government through the project No. PIBA\2023\1\0007(STRAINER). Z.Z. acknowledges support funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034431 and from the “Severo Ochoa" Programme for Centres of Excellence in R&D (CEX2020-001039-S / AEI / 10.13039/501100011033).

### v2

@@ IV.2 Strained moiré patterns
As discussed in previous sections, the strain is a powerful tuning knob for manipulating the size and symmetry of the moiré patterns, much beyond what is possible via only a twist angle. In this section, we provide experimentally examples of nonhexagonal moiré geometries arising due to combinations of twist and strain. Here, the strain could be intentionally generated through well-developed strain techniques (Section IV.1), unintentionally induced during the sample fabrication, or appear due to the lattice relaxation. We will first illustrate a continuous approach to fine tune the twist and strain [55], and then show particular observations of rectangular patterns [88], quasi-one-dimensional patterns [1, 7, 80, 101], and giant atomic swirl deformations [82].
@@ IV.2.1 Continuous tuning moiré patterns
In the moiré systems, the geometric and electronic properties are often highly dependent on the moiré periodicity and the interlayer atomic registry. For example, the superconductivity in TBG is observed around a twist angle 1.1, but has a tolerance of only 0.1 [17]. Therefore, a further optimization of the moiré electronic properties requires a technique to precisely and continuously tune the moiré pattern with twist angle and strain [55, 89].
Kapfer et al. develop a continuous approach to fine tune the twist angle and strain in a graphene–hBN heterostructure, based on the in-plane bending of monolayer ribbons through an AFM tip [55]. Figure 13 illustrates the continuous tuning approach: a monolayer graphene (MLG) ribbon is bent by a few-layer graphite (FLG) manipulator shaped into a gear-like geometry and an AFM tip. The strain induced by bending the ribbon varies both in the longitudinal (x) and transverse (y) directions. The strain map in the bent ribbon follows some general tendencies: (i)  the strain changes linearly in both directions and has the maximum value at the end of the ribbon; (ii)  in the transverse direction, the ribbon has compressive and tensile strain on the inside and outside radius, respectively, which are separated by an unstrained neutral axis; (iii)  the maximum strain in the transverse and longitudinal directions are linearly dependent on the width $W$ and length $L$ of the ribbon, respectively. Therefore, this continuous tuning approach allows one to manipulate independently the twist angle and strain in 2D materials. Importantly, the bending process is robust and reversible.
@@ IV.2.2 Rectangular moiré patterns
Figure 14 shows a strain-induced moiré pattern manipulation in twisted bilayer WSe2 [88]. In this setup, a continuous and homogeneous uniaxial heterostrain is generated by bending a flexible substrate with different radius $R$ (see Sec. IV.1.1). The deformed moiré patterns were visualized by a piezoresponse force microscopy (PFM). As discussed in the previous subsection, in the bending substrate technique the strain is only transferred from the substrate to the bottom layer WSe2, generating a heterostrain in the moiré pattern. In the experiment, the substrate was mounted on several molds with different curvatures, which controlled the strain magnitude from 0% (flat) to 0.5% (radius curvature of 12.5 mm). As seen in Figures 14(b)-(f), this strain variation results in a continuous symmetry modulation of the moiré patterns, from triangular ($\beta_{R}=120^{\circ}$) to rectangular ($\beta_{R}=90^{\circ}$).
Figure 15: Strain induced 1D pattern in TBG. Non-local nano-photocurrent map of a TBG sample with $\theta<0.1^{\circ}$. In the photocurrent map, the bright spots highlight the AA sites. The AB and BA regions are indicated by orange and cyan dots, respectively. In the strained region, double domain walls (DDWs) are formed, separating regions of identical stacking configurations (AB or BA stackings). The green network overlaid on the photocurrent data corresponds to the prediction with one tuning parameter $\bar{\beta}$, indicating the ratio of DDW and single domain wall (SDW) line energies. Adapted under the terms of the $C$ BY license from Ref. [45]. Copyright (2021) The author(s).
@@ IV.2.3 Quasi-1D pattern
As discussed in Section III.4.1, in strained moiré materials there can be a critical strain configuration at which the moiré collapse into a quasi-1D pattern. The critical condition can be, in principle, engineered by extenally induced-strain (Section IV.1). But quasi-1D pattern can also emerge most naturally by unintensional strains, either due to lattice relaxation or sample fabrication. For instance, Figure 15 shows clear rich features of quasi-1D moiré strips of relaxed DWs, for example, the shapes of domains, single domain walls (SDWs), and formation of double domain walls (DDWs) [45]. A dimensionless parameter as the ratio of DDW and SDW line energies was defined to adress a competition between SDWs and DDWs. For $\bar{\beta}<2$, two SDWs attracted each other favoring the emergence of DDW segments. The SDWs and DDWs are also observed in some other moiré materials [45, 1, 46].
Similar strain-induced quasi-1D patterns have been also observed in, for example, TBG strained by a substrate [80], large-angle twisted bilayer WTe2 [125], and exfoliated highly oriented pyrolytic graphite [11].
@@ IV.2.4 Giant atomic swirl
Figure 16: Strain induced giant swirl in bilayer graphene. STM current image ($V_{b}=-300$ mV, $I_{t}=250$ pA) for an intercalated surface of bilayer graphene on SiC, with a deformed moiré showing an (anti-) clockwise swriling feature in (blue) orange. NI indicate the non-intercalated regions. The inset is the STM image of the white dotted square region measued at $V_{b}=900$ mV, $I_{t}=250$ pA. The scal bar is 100 nm. Adapted under the terms of the $C$ BY license from Ref. [82]. Copyright (2023) The author(s).
As noted in Section III.6, relaxation-induced giant atomic swirls can emerge in the domain wall networks of moiré materials [82, 105]. Figure 16 shows a moiré pattern formed in bilayer graphene with only biaxial heterostrain, which was produced by hydrogen intercalation of graphene on SiC [82]. The strain was only imposed to the bottom graphene layer by the SiC reconstruction. Due to the presence of lattice reconstruction in the moiré pattern, domain walls connecting AA stacking formed a triangular network of AB/BA stackings. Notably, giant atomic swirls appeared spontaneously around the AA regions. The neighboring twirl chiralities tend to anti-align (blue and orange arrows). Similar flower-shaped domain walls were already reported in trilayer epitaxial graphene on silicon carbide [64].
@@ VI Acknowledgements
We acknowledge support from NOVMOMAT, Grant PID2022-142162NB-I00 funded by MCIN/AEI/ 10.13039/501100011033 and, by “ERDF A way of making Europe”, and from the ‘Severo Ochoa’ Programme for Centres of Excellence in R&D (CEX2020-001039-S/AEI/10.13039/501100011033). F.E. acknowledges support funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101210351. F.G acknowledges the support from the Department of Education of the Basque Government through the project No. PIBA\2023\1\0007(STRAINER). Z.Z. acknowledges support funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No 101034431.